what are the roots of y=x^2-x-2
Answer by ewatrrr(10682)

 
Hi,
what are the roots of: y=x^2-x-2
x^2-x-2 = 0
Factoring***
(x-2)(x+1) = 0
(x-2) = 0 , x = 2
(x+1) = 0 , x = -1
Checking our factoring using the FOIL method***
F First terms x^2
O Outside terms x
I Inside terms -2x
L Last terms -2